Status of the Top Quark Search

Pekka K. Sinervo University of Toronto / IPP (representing the CDF Collaboration)

1. Why Do We Need to Find Top?

2. The CDF Experiment

3. Search for Dilepton Events

4. Search for Lepton+Jet Events 5. Top Quark Mass Measurement 6. D0 Results

7. Conclusions

CAP Congress

Regina, Saskatchewan June 1994

Why Do We Need a Top Quark

The Standard Model of Elementary Particle Physics

• Three “generations” of quarks and leptons

( \begin{matrix}u \\ d \end{matrix} ) ( \begin{matrix}c \\ s \end{matrix} ) ( \begin{matrix}t \\ b \end{matrix} )
( \begin{matrix}e \\ ν_{e} \end{matrix} ) ( \begin{matrix}μ \\ ν_{μ} \end{matrix} ) ( \begin{matrix}τ \\ ν_{τ} \end{matrix} )

“Isospin” doublets that dictate particle interactions

• Forces carried by bosons

Electromagnetic force  γ
Strong force  eight gluons
Weak force  W ±, Z∘ bosons
Gravity  gravitons

What Happens if Top Doesn’t Exist?

If top doesn’t exist, the SM is in deep trouble

• Masses of the W ± and Z∘ bosons cannot be explained • Properties of the bottom quark imply that its partner exists

– Weak force predicts asymmetry in e+ e− → bb̄ – Not easy to modify SM to accomodate this

• b quark “mixing” could not occur

Properties that a SM Top Quark Must Have

Searches at e^{+} e^{−} Colliders have ruled out “light” top

• data from LEP requires Mtop > 46 GeV/c2 at 95% CL

• Independent of how top decays

LEP measurements of the Z boson properties require

• Mtop = 166{+17+19 \atop −19−22} GeV/c2

Width of W^{±} boson rules out W^{+} → tb̄

• Mtop > 62 GeV/c2 at 95% CL

Searches at pp̄ collisions at CERN and Fermilab

• searched for a “Standard Model” top

• Mtop > 131 GeV/c2 at 95% CL (D0 Collab).

Searching for Top

Top can be pair-produced in proton-antiproton collisions

• Require high collision energies

• Need datasets with high sensitivity

pp̄ Collider at Fermilab is the “only” choice

• High luminosities (L ∼ 5 × 1030 cm−2s−1 )

• Highest energies (√s = 1800 GeV)

Top Quark Signatures

SM top quarks can only decay via the weak force

t → W + b

|_{⟶}ud̄ or cs̄

|_{⟶}e+ νe or μ+ νμ or τ+ ντ

This results in two, potential signatures

• Dilepton events with two, energetic e± and μ∓

• Lepton + Jet events with at least one b quark

The CDF Experiment

CDF is detector designed to study p p interactions

• Collisions with √s = 1.8 TeV created by Tevatron Collider • Large 5000 ton detector

Collaboration consists of

• 35 institutions from US, Japan, Italy, Canada and Taiwan • 440 physicists

The Canadian IPP Group consists of

McGill University  University of Toronto
Faculty  Ken Ragan  Pekka Sinervo
RA  Kurt Biery  Andreas Hölscher

Klaus Strahl

Grad Students Yunling Ye George Sganos
Kostas Kordas  Andreas Warburton

Hyunsoo Kim Wendy Taylor

1992-93 Collider Run

Took data at the Tevatron Collider in 1992-93

• scanned ∼ 1 × 1012 collisions

• Collected 16 × 106 events on tape

• Integrated luminosity of 19.3 pb−1

Most sensitive dataset acquired to date

• If top exists with a mass of 160 GeV/c2 – cross section of 8.2 pb ⇒ 160 produced t t events – expect 1.3 observed dilepton events

– expect 3.5 observed lepton + jet events

Dilepton Top Quark Search

t t decaying into 2 leptons create unique signature

• Two charged leptons from W ± decay

• Imbalance in total energy due to undetected neutrinos • Two jets from bottom quarks

Performed a search requiring

• Two clean lepton candidates with PT > 20 GeV/c • Missing transverse energy E̸T > 25 GeV

• Require two jets with ET > 15 GeV and |η| < 2.0 • Reject specific backgrounds (e.g. Z∘)

Significant dilepton background is Drell-Yan production

• E̸T signature created by mismeasured jet

• Harden E̸T cut if correlated with jet

Backgrounds in Dilepton Channel

Have measured the rate of events from non-top sources

• Have used data to measure these rates

• Cross-checked using Z∘ → e+ e− and Z∘ → μ+ μ−

• Consistent with theoretical calculations

Background  eμ  ee, μμ
Z → τ+ τ−  0.07 ± 0.02  0.06 ± 0.02
One or both leptons fake  0.03 ± 0.03  0.04 ± 0.03
bb̄  0.04 ± 0.03  0.05 ± 0.03
W W, W Z and ZZ  0.10 ± 0.04  0.06 ± 0.02
Drell-Yan  0.10{+0.23 \atop −0.08}
Total  0.24 ± 0.06  0.31{+0.24 \atop −0.10}

Expect 0.56_{−0.13}^{+0.25} background events Observe two events in sample

• Both events have electron and muon candidates

• One has signature of b quark jet

On the other hand, we expect from top production

Mtop (GeV/c2) Number of events
120  3.7
140  2.2
160  1.3
180  0.7

Probability 0.56 events fluctuates to ≥ 2 events is 12%

⇒ On its own, not statistically significant

Mass Limit from Dilepton Data

Can use dilepton data to exclude low mass top

• In this case, remove jet requirement – Jet efficiency less certain for Mtop ∼ 100 − 120 GeV/c2 – No additional events are added to sample

• Add 3.7 pb−1 from 1988-89 data run (no events)

• Backgrounds increase to 3.0 ± 0.6 events

• Assume 2 events are top and set limit on σtop

Lepton + Jets + B Tag Search

Decay mode with the greatest sensitivity is “lepton+jets”

• About 30% of t t events decay this way

• Look for one high PT lepton

• Require E̸T > 25 GeV

• Have 11,900 eνe and 7,000 μνμ candidates

• Require 3 or more jets (from other W decay and b quarks) – ET > 15 GeV and |η| < 2.0

– Find 52 events with these criteria

W + Jet Sample

We have estimated various background sources:

Background  Number of Events
Fake leptons and bb̄  5.2 ± 2.6
W W, W Z and ZZ  2.6 ± 1.2
Z → ττ  1.7 ± 0.5
Z → μμ  1.5 ± 0.4
Z → ee  1.2 ± 0.6
Total  12.2 ± 3.1

Would expect from t t and QCD W+jets

Mtop (GeV/c2) Number of Events
120  39 ± 6
140  22 ± 3
160  12 ± 2
180  6.6 ± 0.7
W+jets  50{+16 \atop −8}

The W+jets remain single largest background

• Reject this by searching for b quark tag

B Tagging Techniques

We use two properties of b quarks to identify them 1. “Long” b lifetime – cτ ∼ 450 microns

• Use the Silicon Vertex Detector to identify jets with a secondary vertex

• We call these “SVX” tags

• Expect 20% of t t events to have at least one SVX tag

2. About 20% of b’s decay via

b → e^{−}ν_{e}X and b → μ^{−}ν_{μ}X

• Look for jets with “soft” leptons in them

• Call these soft-lepton tags (SLT’s)

• Expect about 15% of t t events to have at least one SLT tag

A heavy top would produce fairly energetic b quarks

• For example, for Mtop = 160 GeV/c2 – Have fairly stiff jets

– b quark flight path would be large

SVX B Tags

SVX tag requirements are

• at least 2 tracks with impact parameter d > 3σ

• track PT > 2 GeV/c

• the tracks are consistent with coming from a common vertex

Used control samples to measure the tagging efficiency and fake rate

• inclusive electrons enriched in b → e+ νeX decays

• B → J/ψX, J/ψK±, J/ψK∗ decays

• “generic” jet sample

Inclusive electrons

Have independently estimated 37±8% of sample comes from B decays

• Have measured the rate of single tags and double tags

• Compared +ve pathlength tags with −ve pathlength decays

SVX Tag Backgrounds

The SVX tags found in “generic” jets are due to both real heavy flavour and “mistags”

• The −ve tags come only from “fake” tags

• The +ve tags come from both real c and b quarks and fakes – Use +ve rate to estimate backgrounds to tagged b quarks – results in an overestimate of the background

• Measured tagging rates vs b-jet PT

Tagging Background Estimates

Use fake rates to estimate backgrounds in control samples

• Compare predicted and observed SVX rates in “generic” jet sample

Soft Lepton Tagging

Have good efficiency for identifying muon or electron from b hadron decay

• Require lepton PT > 2 GeV/c

• Require electron to satisfy standard cuts – Use dEdx measurement in CTC to reject pions

– Require muon candidate to have good quality muon stub

Studied SLT tagging using

• photon conversions for electrons (γ → e+ e−)

• J/ψ → μ+ μ− decays

• generic jets to obtain fake rates

Measured efficiency and fake rates

• Efficiency is 0.152 ± 0.023 for Mtop = 160 GeV/c2

• “Fake” tagging rate is ∼ 0.005 − 0.010 tags/track

Have estimated the “fake” rates using various control samples

• Generic jets

• Photon + jet and Z∘ + jet data samples

Fake rates for muons are largest, but still manageable

b Quark Tagging in W+Jet Events

Applied these tagging techniques to W+jet events

• Used fake SVX and SLT tagging rates to estimate backgrounds assuming sample has no b quark content

• The W + 1 and W + 2 jet samples are ultimate control – Don’t expect significant top contribution

– Theory predicts low b hadron content

Sample  SVX bkg  SVX tags  SLT bkg  SLT tags
W+1 jet  15.8 ± 2.1  8  34.1 ± 3.3  33
W+2 jet  6.3 ± 0.8  8  9.9 ± 1.0  12
W+≥ 3 jet  2.30 ± 0.29  6  3.1 ± 0.3  7

Alternative Background Calculation

SVX tag background dominated by “physical” sources

• Background is conservatively overestimated

• Measure fake rate independently using −ve decay lengths • Predict physical backgrounds using theory

Gives a smaller overall SVX background

Sample  SVX bkg (I)  SVX bkg (II)  SVX tags
W+1 jet  15.8 ± 2.1  10.6 ± 3.7  8
W+2 jet  6.3 ± 0.8  4.3 ± 1.4  8
W+≥ 3 jet  2.30 ± 0.29  1.44 ± 0.54  6

Summary of Jet Counting Experiment

We combine the SLT and SVX tagged data

• Compare with background estimates

• Observe 13 tags when we expect 5.4 ± 0.4 background

Features of Tagged Samples

Find that 3 events are “double tagged”

• 2 of these 3 have 2 different jets tagged

• Probability of 2 fake tags is very low

• Expect only 1.3 double-tagged events from background

One dilepton event has jet tagged as an SLT and SVX

• Expect only 0.13 events from purely background

• Expect 0.7 double tags if both events are t t

Statistical Significance

Observe 12 events in total

• 2 dilepton events when we expect 0.56{+0.25 \atop −0.13}

• 13 tags when we expect 5.4 ± 0.4

• 3 double-tagged events

Evaluated the probability of this assuming no top

• Estimate the number of tags from the background sources • Simulate our experiment assuming no top

• Count how often we would see at least the number of tags

Sample  Probability
6 SVX tags  0.032
7 SLT tags  0.038
13 tags + 2 Dilepton events  0.0026

If we use the theoretical prediction for the W +bb̄ background

• Overall significance is 2.6 × 10−4 estimate

Kinematic Features of Events

We have also looked at various kinematic features of events

• Are there other features of data that indicate top

• Good example are energies of second and third jet

• Need to trust MC calculations for W+jet and t t production

Kinematics suggest that of the 52 W + ≥ 3 jet events

• ∼ 20 events from t t production

• remainder mostly QCD W + jets

• Confirmed by looking at ET 2 versus ET 3 for tagged sample

Cross Section for Top Production

Can use our data to estimate top quark cross section

• Subtract the expected backgrounds – Have to “bootstrap” using the tagging efficiencies

– 1.6 ± 0.7 and 1.5 ± 0.7 SVX and SLT background tags • Perform a maximum likelihood fit to estimate σt t

Measurement of Top Mass

We can use the event kinematics to estimate top quark mass

• Require events to have a fourth jet

• Combine all possible jet pairs to form W → qq̄

• Use b-tag information to identify b quark jet

• Fit events requiring kinematic constraints

This selection leaves us with 7 tagged events

• Background estimate is 1.4{+2.0 \atop −1.1} events – Primarily QCD W+jet

Procedure is dependent on theory predictions

Result of Mass Fits

The fitted mass for the 7 events is consistent with t t Fit the mass distribution to obtain

M_{top} = 174 ± 10_{−12}^{+13} GeV/c^{2}

Systematic error is dominated by uncertainties in

• Gluon radiation in t t production (4.4 GeV/c2)

• Different background shapes (∼ 5 GeV/c2)

Log Likelihood Plot

D0 Results

The D0 experiment has also searched in their data

• Looked in di-lepton channels with similar cuts

– Leptons with PT > 12 and 15 GeV/c E̸T

– > 20 GeV

– At least two jets with ET > 15 GeV

• Observe one eμ event – Expect background of 0.98 ± 0.13 events

• Looked at lepton + jet events – Observe 1 electron + jets event

– Observe 2 muon + jet events

– Observe 3 electron + jets + SLT events – Expect background of 3.7 ± 1.0 events

They clearly do not see as large a signal as CDF

• Results are not inconsistent

• Need more data to resolve issue

Are There Other Hypotheses?

We have considered a number of alternative hypothese

1. Are we seeing a b′?

• Cross section would be too large

2. Are we seeing the Higgs boson?

• Only expect to see ∼ 0.2 events at best

3. Cannot rule out exotic scenarios (eg. H+ ) • Will simply need more data

“Best” interpretation is that we are seeing SM top

So Is This a Top Discovery?

Statistically marginal signal

• probability of background fluctuation 0.26%

• Really would like to have ×2 more data

Two consistency checks raise some concerns

1. Predicted rate of QCD W+4 jet events too large

Process  Number of Observed Events
Predicted QCD W+4 jet rate  8 ± 3{+6 \atop −2}
Estimated rate from t t  11.6{+5.6 \atop −4.5}
Observed W+4 jet rate  9

2. Observe an apparent excess of Z + 4 jet events

• Observe 3 Z∘ + 3 jet events and 2 Z∘ + 4 jet events

• Both Z∘ + 4 jet events have b tags!

– Expect 0.64 ± 0.06 background tags

Both observations are statistics-limited

• Cannot rule out statistical fluctuation

• But cannot rationalise these two observations away either

But More Data is Coming!

Tevatron began generating collisions in January 1994

• Run Ib will continue through June 1995

• Double maximum luminosity to 1.5 × 1031 cm−2s−1 • Goal is for CDF/D0 to collect ∼ 75 pb−1

This will allow us to confirm the “top” signal

• After 8 pb−1 have a new e−μ+ event

• Expect to have necessary confirmation by end of 1994 • Start studying top decay modes and other properties

Conclusions

In summary:

• CDF has observed a clear signal for top – 2 Dilepton events

– 13 W+jet events with b quark tags

– Standard Model backgrounds are 5.7_{−0.44}^{+0.49} • Best estimate of top quark mass is

M_{top} = 174 ± 10_{−12}^{+13} GeV/c^{2}

However:

• Statistical significance needs to be improved • Theoretical W +jet rate appears too large

• Details of Z+jet production raise questions • Only consistent with D0 data

More data (coming!) is required to establish claim